OpenStudy (anonymous):

MEDALS!!! The function f(x) = 582(7)^x represents the growth of a mosquito population every year in a remote swamp. Troy wants to manipulate the formula to an equivalent form that calculates every 2 months, not every year. Which function is correct for Troy's purposes?

OpenStudy (anonymous):

f(x) = 582(7 to the one sixth power)6x f(x) = 52(7)x f(x) = 582(712)the x over 12 power f(x) = 5852(7)x

OpenStudy (anonymous):

f(x) = 582(7 to the one sixth power)^6x f(x) = 52(7)^x f(x) = 582(7^12)^x/12 f(x) = 5852(7)^x

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

@sourwing help again!

OpenStudy (anonymous):

anyone!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

f(x) = 582(7 1/6)12x

OpenStudy (anonymous):

THAN YOU SO MUCH! Can you help me with another?

OpenStudy (anonymous):

Tag me in it c:

OpenStudy (anonymous):

Wait thats not an answer choice!

OpenStudy (anonymous):

ugh let me resolve

OpenStudy (anonymous):

can you post a screenshot of the problem?

OpenStudy (anonymous):

I still think that A would work... I'm not 100% sure, though :-/

OpenStudy (anonymous):

uh oh haha ok Ill go with that! Is there a way to test each one?

OpenStudy (anonymous):

Not really :( That problem is such a silly one, it makes no sense...what does 582 have to do with a year or whatever? You're not adding it (so it's not the original amount), you're multiplying it. Sorry I couldn't help with that one... here's a graph for your other problem :)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

No problem c: